The ages of Mark and Adam add up to 28 years total. Mark is 20 years older than Adam. How many years old is Adam?
Found 2 solutions by crazy_mathfreak, lwsshak3:
Answer by crazy_mathfreak(2)
Answer by lwsshak3(11628)

The ages of Mark and Adam add up to 28 years total. Mark is 20 years older than Adam. How many years old is Adam?
**
let x=Adam's age
28-x=Mark's age
x+20=28-x
2x=8
x=4
28-x=24
Adam's age=4
Mark's age=24